eswopes17:

Going against the current, a boat takes 6 hours to make a 120-mile trip. When the boat travels with the current on the return trip, it takes 5 hours. If x = the rate of the boat in still water and y = the rate of the current, which of the following systems could be used to solve the problem?

umm:

'Let speed of boat in still water = x mph & speed of current or stream = y mph Downstream speed = x + y Upstream speed = x - y Using the relation , distance = speed x time, we get, 3(x + y) = 24 or, x + y =8 4(x - y) = 16 or, x - y =4 Solving the system of equations simultaneously, we get, x = 6, y= 2 Rate of current = 2 mph, rate of boat in still water = 6 mph' ~Sources~ https://www.wyzant.com/resources/answers/78697/finding_current_and_rate_of_boat_in_still_water

sillybilly123:

You want: \(5(x + y) = 120\) and \(6(x - y) = 120\)